# WCCNet: Wavelet-context Cooperative Network for Efficient Multispectral Pedestrian Detection

### v2

@@ IV-C Comparison with State-of-the-arts
MMTOD-CG[48] 110 61.40 63.31 70.63 50.26
MMTOD-UNIT[48] 61.54 64.47 70.72 49.43
ThermalDet[49] >40.5>40.5 74.60 78.24 85.52 60.04
ODSC[50] >52.6>52.6 69.62 71.01 82.33 55.53
CFR[42] >16.7>16.7 72.39 74.49 84.91 55.77
GAFF[51] 10.85 72.90 - - -
BU-ATT[52] - 73.10 76.10 87.00 56.10
BU-LTT[52] - 73.20 75.60 86.50 57.40
CMPD[15] 110 69.35 69.64 78,11 59.87
YOLO-MSLite[7] 8.9 75.0 81.70 89.10 54.20
WCCNet 10.50 77.3 84.33 89.51 57.99
WCCNet-L 20.92 78.4 86.16 90.73 57.85
WCCNet-S 6.68 74.7 81.71 88.28 54.06
WCCNet-XS 4.36 70.1 79.94 86.65 43.84
@@
In order to illustrate the versatility of WCCNet when handling other categories, further evaluation is conducted on FLIR which contains multiple types of objects. The experimental results are shown in Table II. Precision results are measured using AP50 for each class and mAP50 over all classes with an IoU threshold of 0.5, which remains the same as in previous works. Table II shows that WCCNet-L gets the best result with $78.4\%$ mAP for overall mAP50, outperforming previous works by at least $4.5\%$. The results demonstrate that our WCCNet is able to generalize well to objects of different categories, not only pedestrians.
IV-D Ablation Study & Further Analysis
@@
In this subsection, we perform ablation experiments on KAIST dataset for detailed quantitative and qualitative analysis of the proposed WCCNet detector. The training settings and other unmentioned hyperparameters are the same across all variants of WCCNet for fair comparison.
Analysis of Dual-stream Multi-domain Backbone
TABLE III: Accuracy and efficiency evaluation of different backbone combinations applied in RGB/Infrared branch.
The conditions with block-out input modality are also evaluated, where ✗  means the block-out layers or modalities.
@@
ADWT ADWT CMRF 10.00 48.92 4.89 26.56 30.66 17.27 11.73 14.12 6.84
ADWT CNN CMRF 17.50 89.98 10.53 21.37 24.56 14.02 9.08 11.44 4.93
CNN ADWT CMRF 17.50 89.98 10.44 19.05 21.17 13.36 7.37 8.55 4.68
@@
ADWT 9.77 46.01 4.36 36.00 42.11 21.18 17.57 21.54 9.12
RGB in Daytime
(a) Origin Image
(b) After DWT Transform
(c) CNN at 2nd Stage
(d) ADWT at 2nd Stage
(e) CNN at 3rd Stage
(f) ADWT at 3rd Stage
RGB at Night
infrared in Daytime
infrared at Night
Figure 6: Origin images, DWT transform results, and feature maps of different branches in WCCNet are shown in different columns, which are grouped by different times including daytime and night. The RGB modality is shown in the top row, while the ones of infrared modality is shown in the bottom row. DWT transform results are simply obtained by the weighted sum of $\mathbf{DH}$ and $\mathbf{DV}$ for schematic representation.
Different Backbone Types. To confirm the efficiency of the proposed dual-stream multi-domain backbone, which extracts infrared features by ADWT and RGB features by CNN, the evaluation on the performance of different backbone combinations has been carried out and summarized in Table III. Meanwhile, feature maps corresponding to these combinations are quantitatively visualized in Fig. 6. Specifically, the backbone with dual CNN subnetworks in the first row of Table III simply replaces the origin ADWT subnetwork in infrared branch with CNN subnetwork. For the backbone with dual ADWT subnetworks in the second row of Table III, two symmetrical CE layers are applied both for the two ADWT subnetworks in CMRF module. The original backbone with paired CNN-ADWT subnetworks is shown in the gray row.
@@
Table III reveals an apparent 18.818.8% increase in precision when substituting backbone with paired ADWT-CNN subnetworks of WCCNet by backbone with paired CNN-ADWT subnetworks. Moreover, there are a noteworthy 35.635.6% reduction in inference time and a 7.47.4% increase in accuracy when backbone with dual CNN subnetworks is replaced with backbone with paired CNN-ADWT subnetworks. Therefore, ADWT is more efficient to extract infrared features than CNN. It is also evidenced in Fig. 6 that, compared with the infrared features extracted by CNN subnetwork, the ones extracted by ADWT subnetwork exhibit more distinct and clearer responses to targets both in daytime and at night. The main reason lays in the fact that there are rich contour features inherent in high-frequency components of infrared images which are perfectly suited for ADWT subnetwork to extract. However, a significant decline in accuracy occurs when dual ADWT subnetworks are applied not only to infrared images but also RGB images, which is further supported by the top two rows in Fig. 6 that CNN subnetwork extracted more distinguished features for RGB modality than ADWT. The above observation shows again that ADWT subnetwork owns stronger feature representation ability to extract contour features from infrared images, while CNN subnetwork performs better at extracting rich textures from RGB images. Consequently, it demonstrates the superiority of our dual-stream multi-domain backbone, which differentially extract features in different domain.
Absence of Different Spectrum. Furthermore, we consider the scenarios where a spectrum is blocked out. In such cases, the corresponding branch for the blocked-out spectrum is simply removed, while CMRF modules are also removed except that CE layers is kept for ADWT. As expected, it can be observed in the last two row of Table III that the accuracy decreases significantly regardless of which modality is removed, and the accuracy of infrared modality is higher than that of RGB modality. The possible reason might be that infrared images are less affected by varying illumination and complex weather for pedestrian targets with thermal radiation, which highlights the essentiality of effective feature extraction for infrared modality.
Analysis of ADWT for Infrared Modality
We delve into ADWT for infrared feature extraction in this subsection, conducting analysis on the impacts of different frequency components and the adaptive mechanism of ADWT.
To understand the importance of different frequency components for infrared modality, four combinations of frequency components extracted by ADWT are evaluated. These four combinations are $\{\mathbf{A}^{j}\}$, $\{\mathbf{DH}^{j},\mathbf{DV}^{j}\}$, $\{\mathbf{A}^{j},\mathbf{DH}^{j},\mathbf{DV}^{j}\}$, and $\{\mathbf{A}^{j},\mathbf{DH}^{j},\mathbf{DV}^{j},\mathbf{HH}^{j}\}$ at $j^{th}$ stage respectively, which represent different frequency subbands and the third one includes the first two ones. Adaptive mechanism is applied for every combination for consistent comparison. The detection precision of different combinations is shown in the top six rows of Table IV, The accuracy of using $\{\mathbf{A}^{j}\}$ alone and using $\{\mathbf{DH}^{j},\mathbf{DV}^{j}\}$ alone is similar, indicating that $\{\mathbf{A}^{j}\}$ and $\{\mathbf{DH}^{j},\mathbf{DV}^{j}\}$ subbands both are indispensable. Besides, when HH components are reserved by ADWT layer, there is a $2.2\%$ reduction in accuracy since HH components are vulnerable to the influence of noises. Consequently, the combination of $\{\mathbf{A}^{j},\mathbf{DH}^{j},\mathbf{DV}^{j}\}$ is chosen to be extracted by ADWT, which outperforms other combinations in precision.
Then we evaluate the detection performance with/without the adaptive mechanism with learnable gain factors $\boldsymbol{\lambda}$. The adaptive mechanism of ADWT applies different weights to different frequency components, which can be regarded as weight function with learnable parameters $\boldsymbol{\lambda}$. We compare vanilla direct concatenation and adaptive mechanism in Table V. The results show that accuracy decay occurs when adaptive mechanism is replaced by direct concatenation.
TABLE IV: Accuracy comparison on different combinations of different infrared frequency components extracted by ADWT.
Frequency Components $\mathrm{MR}^{-2}$ (All) $\mathrm{MR}^{-2}$ (Reasonable)
$\mathbf{A}$ $\mathbf{DH}$ $\mathbf{DV}$ $\mathbf{HH}$ All Day Night All Day Night
19.30 21.80 13.76 7.76 9.24 5.05
19.66 22.33 14.30 8.02 9.42 5.39
19.05 21.17 13.36 7.37 8.55 4.68
19.18 21.96 13.39 7.53 8.75 4.39
TABLE V: Accuracy comparison on using or not using adaptive mechanisms with learnable gain factor $\boldsymbol{\lambda}^{j}$ for ADWT at $j^{t}h$ stage.
Adaptive Mechanism $\mathrm{MR}^{-2}$ (All)
All Day Night
w/o 19.21 21.98 13.42
w/ 19.05 21.17 13.36
Effectiveness of Wavelet-integrated Backbone
To evaluate the efficiency of wavelet-integrated backbone, we proceed with further experiments on the performance of backbones with ADWT applied in different stages. Specifically, the variants of wavelet-integrated backbone gradually substitutes CNN layers by ADWT layers from stage 1 to stage 5. The precision and computational complexity of these variants are evaluated on multiple classification datasets, namely MNIST[53], FashionMNIST[54], and CIFAR-10[55]. The experimental results are shown in Table VI. As ADWT layers gradually replace CNN layers from shallow layers to deep layers, the numbers of parameters and floating point operations progressively decrease, along with the precision.
Grayscale Images. The classification precision slightly decreases when the backbone varies from the highest to the lowest complexity, $0.89\%$ for MNIST and $5.54\%$ for FashionMNIST. At meanwhile, the reduction in network complexity is substantial, almost $89.9\%$ reduction for FLOPs and $91.2\%$ reduction for parameters. MNIST and FashionMNIST datasets contain grayscale images and thus rich contour features, which will be well extracted by ADWT layers in frequency domain.
RGB Images. As for CIFAR-10 containing RGB images with abundant texture features and colors, the variants of backbone which substitute ADWT layers for CNN layers in deeper network exhibit a $32.08\%$ degradation in precision, but still remain $62.06\%$ precision with a tenfold decrease in computational load. The experimental results on RGB images indicate that the Wavelet-integrated backbone is more suitable to extract grayscale features than RGB features.
TABLE VI: Evaluation on ADWT-integrated backbone variants which interchange CNN layers with ADWT layers at different stages. The accuracy and computational complexity are evaluated on classification benchmarks.
@@
MNIST[53] 0.9964 0.9964 0.995 0.9936 0.9875
FashionMNIST[54] 0.9382 0.9368 0.9329 0.9172 0.8862
CIFAR-10[55] 0.9137 0.9142 0.8885 0.8173 0.6206
@@
Early-Cat
(a) Complex Scenario in Daytime
(b) Complex Scenario at Night
Mid-Cat
DMAF
Inter-MA
Late-Cat
CMRF
Figure 7: Comparison on the detection results of WCCNet with different fusion schemes, which are overlaid on RGB/infrared image pairs (in columns). The corresponding attention maps of the detection head are also drawn in the third column. In difficult scenarios, there are occlusions, interfering targets, and small-scale pedestrians. Blue, red and purple boxes stand for true positives, false positives, and false negatives, respectively.
TABLE VII: Accuracy and efficiency evaluation of different fusion methods, along with ablation study on the proposed components of CMRF.
Fusion Scheme
for WCCNet

### v3

@@ 4.3 Quantitative Metrics
MMTOD-CG[MMTOD2019] 110 61.40 63.31 70.63 50.26
MMTOD-UNIT[MMTOD2019] 61.54 64.47 70.72 49.43
ThermalDet[ThermalDet2019] >40.5>40.5 74.60 78.24 85.52 60.04
ODSC[ODSC2020] >52.6>52.6 69.62 71.01 82.33 55.53
CFR[CFR2020] >16.7>16.7 72.39 74.49 84.91 55.77
GAFF[GAFF2021] 10.85 72.90 - - -
BU-ATT[BU2021] - 73.10 76.10 87.00 56.10
BU-LTT[BU2021] - 73.20 75.60 86.50 57.40
CMPD[CMPD2022] 110 69.35 69.64 78,11 59.87
YOLO-MSLite[YOLO-MSLite] 8.9 75.0 81.70 89.10 54.20
ICAFusion[ICAFusion] 26 79.20 81.60 89.00 66.90
WCCNet 10.63 81.66 88.51 92.06 64.42
WCCNet-S 6.74 79.69 87.19 91.67 60.22
WCCNet-XS 4.39 75.82 82.41 88.09 56.97
Table 2: Evaluation results on FLIR dataset. Red, blue, and green colors denote the first, the second, and the third place respectively. - means not reported by the paper of corresponding methods.
4.4 Comparison with State-of-the-arts
4.4.1 Comparison on KAIST
We evaluate WCCNet and conduct comparisons with other state-of-the-art methods [MSDS2018, ARCNN2019, CIAN2019, MBNet2020, MLPD2021, CMPD2022, GAFF2021, MS-DETR2024, RPN2017Konig, zhang2023TripleNet, TCDET2020, yang2022BAANet, ICAFusion, TFNet2024] under reasonable and all-dataset settings following [KAIST2015], as illustrated in Table 1 and Fig 7. The reasonable setting means that the evaluation is conducted on a subset of testing set which only consists of pedestrians taller than 55 pixels without heavy occlusion, while all-dataset setting evaluates detectors on the whole testing set.
Accuracy Comparison
The experimental results on detection precision are given in terms of $\mathrm{MR}^{-2}$. As can be observed in Table 1, all types of our proposed WCCNet outperforms previous methods under all-dataset and reasonable settings. Besides, the MR-FPPI curves under those two settings are shown in Fig. 7(a) and Fig. 7(d) respectively, which also demonstrates the superiority of WCCNet. Especially for all-dataset setting which is more challenging, WCCNet has taken a lead of $13.71\%$ percent over the best recorded method MS-DETR [MS-DETR2024] with 17.81 $\mathrm{MR}^{-2}$ for all-dataset setting. This demonstrates the strong capability of WCCNet to handle small and occlusion pedestrians. Moreover, for detection in low-illumination conditions, the precision results on the night subset under all-dataset setting show that WCCNet surpasses previous methods by a large margin no less than $21.56\%$. Therefore, it confirms that our MoWE-integrated infrared branch can extract sufficient representations of pedestrians in low-illumination conditions with remarkably fewer parameters.
Efficiency Comparison
For evaluation on model efficiency, the number of parameters, FLOPs, and the inference time are provided in Table. 1. Besides, the Speed vs. $\mathrm{MR}^{-2}$ curves under reasonable setting are shown in Fig. 1. It is obvious that, compared with previous works, WCCNet achieves remarkable fast speed with fewer parameters and lower computational cost while maintaining superior accuracy. Especially, WCCNet-S with only 9.75MM parameters and 53.86GG FLOPs can infer at a speed of $152.44$ frames per second (FPS) while maintain competitive accuracy, which again shows the efficiency of WCCNet and its potential in practical applications.
Figure 9: Visualization of extracted and integrated multispectral features of WCCNet. Origin images, features extracted by different branches, fused features by CMRF, and the head features for prediction are shown above, which are grouped by daytime and night. Limited by space, only features at $1^{st}$ and $3^{rd}$ stage are visualized here. As can be observed, pedestrian features are effectively extracted by two different branches, and are further aggregated and refined by CMRF to yield discriminative head features for accurate prediction.
Visual Comparison
Several detection results for complex scenarios are shown in Fig. 8. It can be observed that WCCNet is able to accurately detect pedestrians in challenging situations, including heavy occlusions, absence of RGB information with low illumination, and small-scale pedestrians.
4.4.2 Comparison on FLIR
@@
In order to illustrate the versatility of WCCNet when handling other categories, further evaluation is conducted on FLIR which contains multiple types of objects. The experimental results are shown in Table 2. Precision results are measured using AP50 for each class and mAP50 over all classes with an IoU threshold of 0.5, which remains the same as in previous works. Table 2 shows that WCCNet gets the best result with $81.66\%$ mAP for overall mAP50, outperforming previous works by at least $9.46\%$. The results demonstrate that our WCCNet is able to generalize well to objects of different categories, not only pedestrians.
4.5 Ablation Study & Further Analysis
@@
In this subsection, we perform ablation experiments on KAIST dataset for detailed analysis of WCCNet. The training settings and other unmentioned hyperparameters are the same across all variants of WCCNet for fair comparison.
4.5.1 Analysis of Wavelet-Cooperative Backbone
Different Backbone Types
To confirm the efficiency of the proposed dual-stream backbone, which extracts infrared features by MoWE and RGB features by general neural layers, the evaluation on the performance of different backbone combinations has been carried out and summarized in Table 3. The backbone with dual CNN subnetworks in the first row of Table 3 simply replaces the origin MoWE subnetwork in infrared branch with CNN subnetwork. For the backbone with dual MoWE subnetworks in the second row of Table 3, two symmetrical CE layers are applied for both subnetworks in CMRF module. The backbone of WCCNet is shown in the gray row.
@@ 4.5.1 Analysis of Wavelet-Cooperative Backbone
MoWE MoWE CMRF 10.00 48.96 5.13 24.44 28.90 16.31 10.59 12.49 7.11
MoWE CNN CMRF 17.50 91.00 10.62 20.07 24.09 12.72 8.55 10.54 4.92
CNN MoWE CMRF 17.50 91.00 10.62 17.81 20.36 12.37 6.72 7.80 4.72
@@
MoWE 9.77 46.03 4.42 29.80 36.45 17.65 14.36 17.69 8.28
Table 3: Accuracy and efficiency evaluation of different backbone combinations applied in RGB/Infrared branch. The conditions with block-out input modality are also evaluated, where ✗  means the block-out modules or modalities.
@@
Table 3 reveals an apparent 12.6912.69% decrease in precision when reversing the paired CNN-MoWE subnetworks of WCCNet. Moreover, there are a noteworthy 37.2037.20% reduction in inference time and a 6.366.36% increase in accuracy when backbone with dual CNN subnetworks is replaced with backbone with paired CNN-MoWE subnetworks. Therefore, MoWE is more efficient to extract infrared features than CNN. It is also evidenced in Fig. 9 that, compared with the infrared features extracted by CNN, the approximation features extracted by MoWE exhibit more distinct and clearer responses to targets both in daytime and at night. However, a significant decline in accuracy occurs when dual-MoWE subnetworks are applied, which is further supported by daytime scenarios in Fig. 9 that CNN subnetwork extracted more distinguished features for RGB modality than MoWE. The above observation shows the great significance to asymmetrically build dual-stream wavelet-cooperative backbone which differentially extracts features in different domain.
MoWE Configurations $\mathrm{MR}^{-2}$ (All Setting) Computational Cost
Routed Expert Number $N_{e}$ Balance Loss $\mathcal{L}_{b}$ Shared Expert All Day Night Params. (M) FLOPs (G)
0 19.21 21.98 13.42 17.490458 89.844787
2 18.65 21.77 12.95 17.492704 (+0.002246) 90.169361 (+0.324574)
4 18.12 21.14 12.60 17.493550 (+0.003092) 90.447597 (+0.602810)
8 18.44 21.32 13.18 17.495210 (+0.004752) 90.946725 (+1.101938)
8 18.63 21.55 13.30 17.495242 (+0.004784) 91.004069 (+1.159282)
8 17.81 20.36 12.37 17.495242 (+0.004784) 91.004069 (+1.159282)
16 17.55 20.65 11.89 17.498650 (+0.008192) 92.117013 (+2.272226)
32 17.43 20.32 10.73 17.505394 (+0.014936) 94.342903 (+4.498116)
Table 4: Accuracy and efficiency evaluation with different expert configurations. Metrics are evaluated on KAIST test set. The activation ratio of routed experts is set to 0.5.
Visualization of Feature Extraction and Cooperation
We visualize extracted features of different branches, fused features by CMRF, and the head features for prediction in Fig. 9. It can be observed that, for night scenarios, pedestrian features extracted by MoWE are more prominent than those extracted by CNN, which CNN achieve better at daytime. Besides, compared with solely extracted features in one branch, features fused by CMRF exhibit more concentrated responses to pedestrians, which demonstrates the effectiveness of CMRF. The final head features show that WCCNet can accurately locate pedestrians for complex scenarios.
Figure 10: Accuracy vs. efficiency analysis of varying routed expert numbers. Detection Accuracy $MR^{-2}$ is compared with model parameters and FLOPs.
4.5.2 Analysis of MoWE
Analysis of Different Expert Configurations
We conduct ablation experiments on three main configurations of MoWE, i.e., routed expert number $N_{e}$, balance loss $\mathcal{L}_{b}$, and shared expert, as shown in Table 4 and Fig. 10. There is an improvement of $7.29\%$ in accuracy when $N_{e}$ change from 0 to 8, demonstrating the significance of routed expert ADWT. As the number of routed experts $N_{e}$ increases from 0 to 32, the accuracy and computational cost simultaneously increase. Fig. 10 shows that $N_{e}$=8 meets a trade-off between accuracy and efficiency. Besides, compared with generic neural layers, the increment of computational overhead brought by MoWE is marginal, i.e., less than 0.08% in paramters and 5.01% in FLOPs, which highlights the efficiency of MoWE.
The involvement of balance loss $\mathcal{L}_{b}$ and shared expert can further improve the accuracy by $3.42\%$ and $4.40\%$ respectively. The balance loss help to prevent routing collapse and fully explore all the available experts, while the shared expert can further enhance the feature extraction stability of MoWE.
Figure 11: Analysis of routed expert activation patterns for MoWE over the entire KAIST test set. (a) compares the activated ratio and corresponding distribution across different experts in left and right sub-figures respectively. (b) illustrates the activation frequency distribution across different channels. (c) shows the activation frequencies of experts across samples for daytime and night scenarios.
Analysis of Expert Activation Patterns
Fig. 11a, Fig. 11b, and Fig. 11c show the activation patterns across different experts, channels, and samples respectively. Fig. 11a shows that the activation ratio of each expert is approximately equal, which demonstrates that the routing mechanism is balanced. Fig. 11b indicates that each channel has their preferred experts. That is, different experts have their own strengths in feature extraction for different channels. It can be observed from Fig. 11c that the activation patterns of experts vary distinctively between daytime and night-time samples. The activation frequencies become more stable across night-time samples, and Expert $\{3,4,7\}$ exhibit higher usage. This behavior is attributed to the fact that night-time infrared images possess more salient pedestrian features, which are easier for MoWE to capture than their daytime counterparts.
Figure 12: Comparison on fused features $\mathbf{G}_{T}^{j}$ with/without CMRF. Fused $\mathbf{G}_{T}^{j}$ features in CMRF are visualized by heatmaps and 3D curves in different columns, which are grouped by different times including daytime and night. When not applying CMRF, $\mathbf{G}_{T}^{j}$ refers to features of Mid-Cat fusion. As can be observed, feature maps of CMRF effectively concentrate on spatial regions associated with pedestrians, suppressing interference from similar backgrounds.
@@ 4.5.2 Analysis of MoWE
MNIST 0.9964 0.9964 0.995 0.9936 0.9875
FashionMNIST 0.9382 0.9368 0.9329 0.9172 0.8862
CIFAR-10 0.9137 0.9142 0.8885 0.8173 0.6206
@@
Table 5: Evaluation on ADWT-integrated backbone variants which interchange CNN layers with ADWT layers at different stages. Grayscale and RGB images of three classification benchmarks are involved.
Effectiveness of ADWT-integrated Neural Network
To evaluate the efficiency of ADWT, we extract RGB branch of WCCNet as classification model, and progressively replaced CNN layers with ADWT layers from stages 1-5. Evaluations are conducted on grayscale datasets MNIST and FashionMNIST, and RGB dataset CIFAR-10, as shown in Table 5. Increasing ADWT substitution significantly reduces complexity, up to $89.9\%$ FLOPs and $91.2\%$ parameters, but trades off precision. Grayscale datasets show minimal accuracy degradation, $0.89\%$ for MNIST and $5.54\%$ for FashionMNIST, benefiting from frequency-domain processing of ADWT. However, RGB data (CIFAR-10) suffered a $32.08\%$ drop in deeper substitutions despite a tenfold efficiency gain, indicating ADWT is more effective for grayscale than RGB feature extraction.
4.5.3 Analysis of Crossmodal Rearranging Fusion
Quantitative Comparisons with Different Fusion Schemes
Table 6 shows ablation studies on comparison between our proposed CMRF and other various fusion schemes. The different fusion schemes are all implemented within WCCNet framework. The Early Concatenation (Early-Cat) and Late Concatenation (Late-Cat) fusion schemes only concatenate multimodality feature maps at the first and the last stage of the backbone along the channel dimension respectively, and get the lowest precision. The Middle Concatenation (Mid-Cat) fusion scheme concatenates RGB and infrared features along the channel dimension at every stage of the backbone, and thus obtain higher accuracy with a minor decline in $MR^{-2}$. These channel-dimension concatenation-based fusion schemes neglect the interaction between multimodal features which is essential for extracting complementary information and filtering out redundant information, hence unsurprisingly they get the lowest precision.
The middle fusion schemes designed for crossmodal interaction including DMAF, Inter-MA, and CMRF, are also evaluated. It is worth mentioning that the features for middle fusion are all extracted after CE layer. The DMAF[MBNet2020] and Inter-MA[zhang2023TripleNet] both obtain complementary information by subtracting feature maps of different modalities along the channel dimension to extract numerically differential features. Instead, the proposed CMRF extracts complementary information through mining semantic-aware features in spatially relevant local regions shared by different modalities.
Among the above fusion schemes, CMRF achieves the highest accuracy with competitive inference time, outperforming the second-best DMAF by $13.71\%$ in accuracy.
Fusion Scheme for WCCNet